OpenStudy (anonymous):

I Don't Understand Quadratic Equation Equations: Solve 3x^2 + 5x + 7 = 0. Round Solutions To The Nearest Hundredth

OpenStudy (anonymous):

3x² - 5x - 7 = 0 3x² - 5x = 7 x² - 5/3x = 7/3 x² - 5/6x = 84/36 + 25/36 (x - 5/6)² = 109/36 x - 5/6 = 1.740051 1st factor: = x - 0.833333 + 1.740051 = x + 0.906718 2nd factor: = (3x² - 5x - 7)/(x + 0.906718) = 3x - 7.720154 Values of x: x + 0.906718 = 0, x = - 0.906718 3x - 7.720154 = 0, 3x = 7.720154, x = 2.573385 Answer: x = - 0.906718, 2.573385 Proof: 3(- 0.906718)² - 5(- 0.906718) = 7 3(0.822138) + 4.53359 = 7 2.46641 + 4.53359 = 7 7 = 7 Proof: 3(2.573385)² - 5(2.573385) = 7 3(6.62231) - 12.8669 = 7 19.8669 - 12.8669 = 7 7 = 7

OpenStudy (anonymous):

You can also just use the quadratic equation formula.  Just set in 3 to A 5 to B and 7 to C